# Scaling Generalist Data-Analytic Agents

### v1

@@ Appendix B Limitations
This paper still has some limitations that must be acknowledged: a) At present, we only incorporate reasoning-oriented data-analysis tasks; training, predictive, and data-visualization tasks are deliberately excluded and reserved as our important future work. b) Owing to computational constraints, our experimental backbone is restricted to the Qwen family, with model scale capped at 14B. Furthermore, not all mainstream benchmarks are covered in our evaluation suite. c) Limited by computational resources, we have not exhaustively evaluated all RL training algorithms; moreover, data scarcity constrains our RL runs to $\sim 350$ steps. In future work, we will investigate more advanced RL strategies that enable stable, continual learning over substantially larger datasets.
@@ Appendix C A More Detailed Related Work
The release of Large Reasoning Models (OpenAI, 2024c; DeepSeek-AI et al., 2025; Kimi et al., 2025) marks the triumph of Reinforcement Learning (RL) for LLMs. GRPO-style algorithms (Shao et al., 2024; Yu et al., 2025; Yue et al., 2025b; Wang et al., 2025; Dong et al., 2025; Zhang et al., 2025b) enable models to autonomously explore while preserving robust generalization across diverse reasoning patterns. Consequently, a surge of work has sought to adapt GRPO-like algorithms to various agent domains (Jin et al., 2025a; Song et al., 2025; Li et al., 2025c; Wu et al., 2025a; Feng et al., 2025; Qian et al., 2025; Li et al., 2025b; Wei et al., 2025b). Yet these methods presuppose a strong backbone model; researchers are therefore compelled to synthesize copious post-training data to compensate for the backbone’s deficiencies in the target agent setting. To the best of our knowledge, we are the first to systematically investigate the scaling of agent post-training data and multi-turn RL training in the data-analytic scenario, aiming to provide actionable insights for data synthesis and RL-driven training in other complex agent fields.
@@ Appendix E Baselines and Reproduction Details
We compare our DataMind with five strong proprietary models: GPT-4o (gpt-4o-2024-0806) (OpenAI, 2024a), o4-mini (o4-mini-2025-04-16) (OpenAI, 2025b), DeepSeek-R1 (deepseek-r1-2025-0528) (DeepSeek-AI et al., 2025), DeepSeek-V3.1 (deepseek-v3.1-nothinking) (DeepSeek-AI, 2025), GPT-5 (gpt-5-2025-08-07) (OpenAI, 2025a). We also include four outstanding open-source models: QwQ-32B (Qwen, 2025), Qwen-2.5-Coder-32B (Hui et al., 2024), Llama-3.3-70B (Dubey et al., 2024), and Qwen-2.5-72B (Yang et al., 2024). In addition, we select four open-source models that have been explicitly trained for data-analysis-related tasks: TableLLM (Wu et al., 2025b) and Table-R1 (Wu et al., 2025c) are specialized for tabular reasoning, whereas OmniSQL (Li et al., 2025a) and SQL-R1 (Ma et al., 2025) are optimized for Text-to-SQL generation. We include Qwen-2.5-Coder-7B and 14B (Hui et al., 2024) as backbone models to compare different baselines. We use the Instruct version for all open-source models. Here we introduce the baselines we compare with and our reproduction details:
@@ Appendix F Training and Inference Details
We use LlamaFactory (Zheng et al., 2024) for SFT training and verl (Sheng et al., 2025) for RL training. For SFT, our learning rate is 1e51e-5 with a warmup ratio of $0.1$ and a cosine decay schedule. Our global batch size is set to $16$. For RL, we use a learning rate of 1e61e-6. The batch size is $16$ with a mini batch size of $2$. The rollout temperature is $0.7$, the top-p is $1.0$, and the group size $G$ is $4$. We schedule $\gamma$ via cosine decay, annealing from a peak of $0.9$ to a valley of $0.05$. At test time, we fix the temperature to $0.7$, top-p to $0.95$, and an inference batch size of 5 for all evaluations. The detailed hyperparameters employed in DataMind are presented in Tab.2.
@@
     batch size      16
     mini batch size      2
@@
Appendix G Prompts Used in Our Paper
G.1 Training and Evaluation Prompt
@@
G.2 Query Synthesis Prompt
@@
G.3 Trajectory Sampling Prompt
@@ 6. Important Notes:
G.4 Judge Model Prompt

### v2

@@ Appendix B Limitations
This paper still has some limitations that must be acknowledged: a) At present, we only incorporate reasoning-oriented data-analysis tasks; training, predictive, and data-visualization tasks are deliberately excluded and reserved as our important future work. b) Owing to computational constraints, our experimental backbone is restricted to the Qwen family, with model scale capped at 14B. Furthermore, not all mainstream benchmarks are covered in our evaluation suite. c) Limited by computational resources, we have not exhaustively evaluated all RL training algorithms; moreover, data scarcity constrains our RL runs to $\sim 350$ steps. In future work, we will investigate more advanced RL strategies that enable stable, continual learning over substantially larger datasets. d) The current version of DataMind only accepts tabular data files and textual questions. In the future, we will extend DataMind to additional modalities.
@@ Appendix C A More Detailed Related Work
The release of Large Reasoning Models (OpenAI, 2024c; DeepSeek-AI et al., 2025; Kimi et al., 2025) marks the triumph of Reinforcement Learning (RL) for LLMs. GRPO-style algorithms (Shao et al., 2024; Yu et al., 2025; Yue et al., 2025b; Wang et al., 2025a; Dong et al., 2025; Zhang et al., 2025b) enable models to autonomously explore while preserving robust generalization across diverse reasoning patterns. Consequently, a surge of work has sought to adapt GRPO-like algorithms to various agent domains (Jin et al., 2025a; Song et al., 2025; Li et al., 2025c; Wu et al., 2025a; Feng et al., 2025; Qian et al., 2025; Li et al., 2025b; Wei et al., 2025b). Yet these methods presuppose a strong backbone model; researchers are therefore compelled to synthesize copious post-training data to compensate for the backbone’s deficiencies in the target agent setting. To the best of our knowledge, we are the first to systematically investigate the scaling of agent post-training data and multi-turn RL training in the data-analytic scenario, aiming to provide actionable insights for data synthesis and RL-driven training in other complex agent fields.
@@ Appendix E Baselines and Reproduction Details
We compare our DataMind with five strong proprietary models: GPT-4o (gpt-4o-2024-0806) (OpenAI, 2024b), o4-mini (o4-mini-2025-04-16) (OpenAI, 2025b), DeepSeek-R1 (deepseek-r1-2025-0528) (DeepSeek-AI et al., 2025), DeepSeek-V3.1 (deepseek-v3.1-nothinking) (DeepSeek-AI, 2025), GPT-5 (gpt-5-2025-08-07) (OpenAI, 2025a). We also include four outstanding open-source models: QwQ-32B (Qwen, 2025), Qwen-2.5-Coder-32B (Hui et al., 2024), Llama-3.3-70B (Dubey et al., 2024), and Qwen-2.5-72B (Yang et al., 2024). In addition, we select four open-source models that have been explicitly trained for data-analysis-related tasks: TableLLM (Wu et al., 2025b) and Table-R1 (Wu et al., 2025c) are specialized for tabular reasoning, whereas OmniSQL (Li et al., 2025a) and SQL-R1 (Ma et al., 2025) are optimized for Text-to-SQL generation. We include Qwen-2.5-Coder-7B and 14B (Hui et al., 2024) as backbone models to compare different baselines. We use the Instruct version for all open-source models. Here we introduce the baselines we compare with and our reproduction details:
@@
Note that for a fair comparison, we strictly adhere to the original input–output formats and benchmark configurations of every baseline. Specifically, for TableLLM and Table-R1, as they are both trained on TableInstruct, we evaluate them with the same prompt used in TableInstruct and converted each table file from all the benchmarks into the string representation required by their task settings before inserting it into the prompt. For OmniSQL and SQL-R1, which are both trained on Text-to-SQL tasks, we follow their native prompts and ask the models to generate complete SQL statements directly based on the table schema and the user query. In addition, we uniformly transform all benchmark data files into the .sqlite format via scripts to match their task requirements.
@@ Appendix F Training and Inference Details
We use LlamaFactory (Zheng et al., 2024) for SFT training and verl (Sheng et al., 2025) for RL training. For SFT, our learning rate is 1e51e-5 with a warmup ratio of $0.1$ and a cosine decay schedule. Our global batch size is set to $16$. For RL, we use a learning rate of 1e61e-6. The batch size is $32$ with a mini batch size of $4$. The rollout temperature is $0.7$, the top-p is $1.0$, and the group size $G$ is $4$. We schedule $\gamma$ via cosine decay, annealing from a peak of $0.9$ to a valley of $0.05$. At test time, we fix the temperature to $0.7$, top-p to $0.95$, and an inference batch size of 5 for all evaluations. For all the processes, the maximum number of interaction rounds $\mathcal{T}$ is set to 10. Each of the training experiments can be run on a machine with 8 80G A100 GPUs within 2 days. The detailed hyperparameters employed in DataMind are presented in Tab.3.
@@
     batch size      32
     mini batch size      4
@@
Appendix G More Results
G.1 More Base Models.
We primarily conduct experiments with the Qwen family, as it is widely acknowledged that Qwen exhibits stronger reasoning capacity and higher plasticity during RL training (Gandhi et al., 2025; Wang et al., 2025b). To verify the universality of our DataMind pipeline on different kinds of base models, we additionally train Llama-3.1-8B-Instruct (Meta, 2024) on DataMind-14K and the results are reported in Tab.4. It can be observed that the DataMind pipeline transfers robustly across model families. Llama-8B not only surpasses Llama-70B but also performs on par with strong baselines. Its pass@3 further approaches the level of DeepSeek-V3.1 and DataMind-7B.
Table 4: Results of Llama-8B Model compared with other baselines.
Backbone Method DABench TableBench BIRD Avg.
pass@1 pass@3 pass@1 pass@3 pass@1 pass@3 pass@1 pass@3
DeepSeek-V3.1 ReAct 81.32 89.49 72.52 81.68 57.89 68.12 70.58 79.76
GPT-5 78.21 85.21 69.93 78.37 60.17 65.19 69.44 76.26
Llama-3.3-70B ReAct 69.78 80.16 55.47 70.36 59.10 68.58 61.45 73.03
Qwen-2.5-72B 75.33 86.38 65.44 76.21 60.30 69.49 67.02 77.36
Qwen-2.5-Coder-7B DataMind 77.30 87.94 67.60 79.39 59.41 69.88 68.10 79.07
Llama-3.1-8B-Instruct DataMind 73.67 87.16 64.49 76.46 58.52 69.56 65.56 77.73
G.2 More Baselines.
Data Interpreter has been embedded into the MetaGPT framework [^3] https://github.com/FoundationAgents/MetaGPT. as an off-the-shelf analytics tool. Despite our best effort to decouple its pipeline and run it in isolation, we are unable to reproduce the reported 94.93% GPT-4o score on DABench and it also cannot be adapted to SQL tasks. The same discrepancy has been independently acknowledged in another paper (You et al., 2025). Our reproduction results are shown in Tab.5. It can be found that task-specific scaffolds do not consistently outperform the vanilla ReAct paradigm, and prompts engineering for one model may fail to generalize to others. We believe that the greatest truths are the simplest. This is also why we believe agent training, rather than prompt engineering and scaffold constructing, is the more promising path.
Table 5: Results of Data Interpreter compared with other baselines.
Backbone Method DABench TableBench
pass@1 pass@3 pass@1 pass@3
DeepSeek-V3.1 ReAct 81.32 89.49 72.52 81.68
Data Interpreter 67.70 80.93 61.98 74.52
Qwen-2.5-72B ReAct 75.33 86.38 65.44 76.21
Data Interpreter 62.26 75.10 49.90 58.84
Qwen-2.5-Coder-7B ReAct 15.05 35.41 11.70 28.63
Data Interpreter 54.09 65.37 27.99 41.82
DataMind 77.30 87.94 67.60 79.39
Qwen-2.5-Coder-14B ReAct 71.21 83.27 56.96 69.97
Data Interpreter 60.70 74.71 40.71 55.82
DataMind 80.29 88.72 70.95 81.81
Table 6: Results on QRData.
Backbone Method QRData
pass@1 pass@3
DeepSeek-V3.1 ReAct 60.75 75.67
Qwen-2.5-72B ReAct 60.50 72.75
Qwen-2.5-7B DataMind 57.66 69.34
Qwen-2.5-14B DataMind 62.04 77.62
G.3 More Benchmarks.
To demonstrate that DataMind can generalize to more challenging data-analysis benchmarks, we further evaluate on QRData (Liu et al., 2024), a dataset that features intricate causal-reasoning tasks and relies heavily on domain commonsense knowledge. We compare our DataMind-7B and DataMind-14B agents against the strongest open-source and closed-source models; results are reported in Tab.6. Our 14B model still achieves the best performance, benefiting from the comprehensive coverage of data-analysis types and domains in DataMind-12K, as well as the consistent performance gains brought by our stable agent training strategy. Our 7B model also delivers competitive results. However, constrained by its parameter count, it lacks the domain-specific knowledge required, which prevents it from achieving equally impressive performance.
G.4 More Evaluation Metrics.
Table 7: Rouge-L Results of various methods compared with our model-as-judge.
Backbone Method TableBench
Rouge-L pass@1
DeepSeek-V3.1 ReAct 19.64 72.52
Qwen-2.5-72B 17.93 65.44
Qwen-2.5-7B TableLLM 11.33 9.81
Table-R1 13.86 25.49
DataMind 23.41 76.47
To further validate the soundness of our model-as-judge protocol, we re-score the descriptive tasks in TableBench with Rouge-L and contrast these scores with the ratings previously assigned by GPT-4o-mini. The comparison is reported in Tab.7. The resulting rankings align closely with those produced by our model-as-judge. Nevertheless, Rouge-L clearly fails to capture true model performance. It over-emphasizes surface lexical and sentence overlap with the gold label rather than answer correctness. This renders the performance differences among models marginal. Even the powerful DeepSeek-V3.1 attains a very low Rouge-L score, underscoring the rationality and fairness of employing a judge model for evaluation.
Table 8: Table Header Overlap Ratio between DataMind and all the benchmarks.
DABench TableBench BIRD
DataMind 1.29% 0.00% 0.02%
Appendix H Data Contamination Analysis
To the best of our knowledge, the DABench data files are harvested from GitHub, whereas the tables in TableBench are extracted from Wikipedia, with neither overlapping with the corpora we curate. We sample the training split of BIRD and adopt synthetically generated schemas provided by OmniSQL. Consequently, our sources are also disjoint from the BIRD test set. To quantify this claim, we compute the header-overlap ratio between every table in DataMind-12K and the headers of each benchmark in Tab.8. The resulting overlap is exactly 0% (or statistically indistinguishable from 0%) for every benchmark, demonstrating that DataMind-12K introduces no data contamination into the evaluation.
Appendix I Prompts Used in Our Paper
I.1 Training and Evaluation Prompt
@@
I.2 Query Synthesis Prompt
@@
I.3 Trajectory Sampling Prompt
@@ 6. Important Notes:
I.4 Judge Model Prompt